OpenStudy (anonymous):

If you have 310.0 mL of water at 25.00 °C and add 140.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

Using the first law of thermodynamics (I'm not sure about if it's the first) ΔU = Q + ΔW=0( you need more hypothesis to do this) ΔU_1+ΔU _2=0 m_1 : mass of 310.0 mL of water m_2 : mass of 310.0 mL \[m_1c(T_f-T_0)+m_2c(T_f-T_1)=0\] \[T_f=\frac{m_1T_1+m_2T_2}{m_1+m_2}\] is that chemistry or thermodynamic !

OpenStudy (anonymous):

uhh..chemistry

OpenStudy (anonymous):

uhhh.. i don't understand the formula